OpenStudy (anonymous):

Find the derivative of f(x) = 5 divided by x at x = -1.

OpenStudy (jdoe0001):

use the power rule, keep in mind that \(\bf f(x)=\cfrac{5}{x}\implies f(x)=5x^{-1}\)

OpenStudy (anonymous):

okay can u show me how to go from there, 5x^-1 is probably easier to use in the power rule formula right?

OpenStudy (jdoe0001):

yes... the other way would be using the product formula, which will be a bit longer, is all

OpenStudy (jdoe0001):

rather the "quotient" formula

OpenStudy (jdoe0001):

\(\bf f(x)=\cfrac{5}{x}\implies f(x)=5x^{{\color{red}{ -1}}}\implies f(x)={\color{red}{ -1}}\cdot 5\cdot x^{{\color{red}{ -1}}-1}\)

OpenStudy (jdoe0001):

that'd give you the derivative equation, then just plug in "-1" for its "x" argument

OpenStudy (anonymous):

okay so would the derivative be -5?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

what is -1-1

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

that is your exponent on x

OpenStudy (anonymous):

okay so x^-2 or 1/x^2 right?

OpenStudy (jdoe0001):

\(\bf f(x)=\cfrac{5}{x}\implies f(x)=5x^{{\color{red}{ -1}}}\implies f(x)={\color{red}{ -1}}\cdot 5\cdot x^{{\color{red}{ -1}}-1} \\ \quad \\ f'(x)=-5x^{-2}\qquad \textit{when x = -1, thus}\quad f'({\color{blue}{ -1}})=-5({\color{blue}{ -1}})^{-2}\)

OpenStudy (anonymous):

ohhhhhh okayy so i would simplify then plug in the x=-1 in for x in the new equation

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

okay so the derivative is -5 right?

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

the derivative isnt -5. but at f(-1)=-5

OpenStudy (anonymous):

f'(-1)

OpenStudy (jdoe0001):

right

OpenStudy (anonymous):

ohhhhhh okayyyy thank youuu so muchhh, really helpfull

OpenStudy (jdoe0001):

yw